A square-shaped copper coil has edges of length 50 cm and contains 50 turns. It is placed perpendicular to a 1.0 T magnetic field. It is removed from the magnetic field in 0.25 s and restored in its original place in the next 0.25 s. Find the magnitude of the average emf induced in the loop during (a) its removal, (b) its restoration and (c) its motion.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. (a) During removal, ϕ base 1 = B.A. = 1 x 50 x 0.5 x 0.5 – 25 x 0.5 = 12.5 Tesla-m^2 ϕ base 2 = 0, t = 0.25 e = - dϕ/dt = ϕ base 2 – ϕ base 1/dt = 12.5 /0.25 = 125x 10^-1 / 25x 10^-2 = 50V (b) During its restoration ϕ base 1 = 0 ; ϕ base 2 = 12.55 Tesla- m^22 ; t = 0.25 s E = 12.5 -0 /0.25 = 50 V. (c) During the motion Φ base 1 = 0, ϕ base 2 = 0 E dϕ/dt = 0